# Asymmetric Influence of the Amplitude-Dependent Tune Shift on the Transverse Mode-Coupling Instability

### v1

@@ IV.2 Beam Losses at Thresholds
A significant influence of the value and sign of the ADTS coefficients can be observed in the amount of charge lost when the instability threshold is crossed during injection. Figure 7 shows the beam loss in percent for different values of the ADTS coefficient with a chromaticity close to zero or with slightly negative chromaticity values. For negative ADTS coefficients up to nearly zero ($\approx-500/\mathrm{m}$ ) no beam loss is encountered at all when crossing the instability threshold during injection. This is already noteworthy as it shows, that the instability is not destructive even though it leads to strong center-of-mass oscillations and an increase in bunch size. On the other side, for positive ADTS coefficients a partial beam loss is observed when crossing the threshold. For values from zero up to $6000/\mathrm{m}$ more than 40 and up to 90 percent of the beam current is lost. For higher positive ADTS coefficients the loss goes down close to zero again. So, for ADTS coefficients up to $6000/\mathrm{m}$ there is a difference in the observed behaviour for a positive and negative sign of the ADTS coefficient. While at negative coefficients the instability is self-containing, for positive coefficients a partial beam loss is observed until the beam stabilizes again.
@@ IV.3 Bunch Position and Size
Figure 8: Synchronized measurement of vertical spot profile and center-of-mass amplitude as function of time. The image shows the vertical light spot measured at different points in the sawtooth like dynamic. In red the envelope of the center-of-mass motion is displayed with a point for each spot profile measurement. The yellow dots indicate the spot size gained from a Gaussian fit. Measured at ADTS coefficient $b=-10000$ and current of 2.6 mA.
For a negative ADTS coefficient the dynamic above threshold shows clear, regular, sawtooth like bursts in bunch size and as amplitude modulation of the center-of-mass oscillations. For measurements this is visible in the BPM trace directly (Fig. 2) as well as in the synchronous measurement of bunch position and bunch size in Fig. 8. The contribution of the bunch size can be seen in the fact that the spot size goes down slower than the center-of-mass oscillations seen by the BPMs. Additionally, the calculated spot size of up to 1.7 mm (determined by Gaussian fit) in Fig. 8 compared to the maximal detected center-of-mass oscillation amplitude of 0.25 mm (= 0.5 mm peak-peak) indicates that the bunch size has a non-negligible contribution. Furthermore, the center-of-mass oscillation goes back to nearly zero for some hundred turns between the increases in oscillation amplitude. During this time the observed spot size slowly damps down indicating that this corresponds to the bunch size damping down, reaching minimal values in the order of 0.3 mm before the next peak. The same behaviour is present in the simulation shown in Fig. 9.
Figure 9: Simulated center-of-mass oscillation and bunch size at an ADTS coefficient $b=-15000$ and a current of 4.8 mA.
@@
This dynamic indicates a stabilizing mechanism which leads to a containment of the instability instead of a continuous growth until charge is lost. The observed behavior in the amplitude of the center-of-mass oscillations and the bunch size show that at some point a temporary stabilization occurs which leads to a damping down of the oscillation to below the noise limit of the measurement. The bunch size is also damped down during this stable period but it does not reach the expected stable bunch size before the instability is triggered again leading to a fast blow up of the bunch size and the onset of strong center-of-mass oscillations. A possible mechanism could be via an increased tune spread due to the blown up bunch size and the ADTS leading to an increased Landau damping effect. While the IBS could be the cause of the hysteresis observed in the Instability threshold, it is not a candidate for explaining the self-containing dynamics, as it operates in the wrong sense, ie. a blown-up beam has a lower threshold current and is therefore more unstable and does not contribute to a self-containing effect, where the threshold would need to increase to temporarily stabilize the unstable beam.
@@
For positive ADTS, the dynamics above the threshold can only be observed in measurements at high ADTS coefficients where no instantaneous charge loss occurs. For higher bunch currents the dynamics in the center-of-mass oscillation and the bunch size have a similar sawtooth like pattern as observed for negative ADTS coefficients (see left side of Fig. 10). The pattern changes for lower bunch currents as shown on the right side in Fig. 10. Here, the center-of-mass oscillation amplitude is more constant and lightly modulated in the measurements, and nearly constant in the simulations.
Comparing the simulated maximal oscillation amplitudes and bunch sizes reached in, for example, Fig. 9 and the lower left plot of Fig. 10, shows a stronger blow up of the beam for the positive sign of the ADTS. The same difference is observed in measurements when comparing the maximal amplitude of the center-of-mass oscillation in Fig. 2 and the upper left plot in Fig. 10.
@@
This indicates, that the level of “blow up” at which the instability is contained and finds some kind of equilibrium, pseudo-stable state is different for negative and positive ADTS. The asymmetry is visible clearly in Fig. 11, where for a bunch current slightly above threshold, the maximal bunch size and the maximal oscillation amplitude of the center-of-mass is given as a function of ADTS coefficient for simulations [^1] To be more robust against outliers the 95th percentile of the bunch size and the center-of-mass oscillation amplitude are taken.. The range in ADTS coefficient where partial current loss would occur probably depends on the combination of the center-of-mass oscillations and the total bunch size which, above a certain value, would lead to parts of the charge being “scraped” by the beam pipe. Independent of the exact value, it can be seen from Fig. 11 that the affected ADTS range would not be symmetric around zero but rather shifted to positive ADTS coefficients. In measurements, the same dependence of the maximal center-of-mass oscillation amplitude [^2] Again, the 95th percentile of center-of-mass oscillation is taken. as a function of the ADTS coefficient is observed (Fig. 12) for negative ADTS. While it is not measurable at the lower positive ADTS coefficients, due to the partial beam losses, the measured values at higher positive ADTS coefficients are higher than the corresponding values at negative ADTS showing the same asymmetry as in the simulations in Fig. 11. So to reach the same level of suppression of the instability, meaning low values in maximal bunch size and center-of-mass oscillations, a higher positive than negative ADTS coefficient would be needed. This asymmetry observed in both measurement and simulations supports the explanation for the asymmetry observed in the beam loss at the threshold (see Fig. 7).
@@
Figure 11: Simulated maximal (95percentile) bunch size and center-of-mass oscillation amplitude and minimal (5percentile) action $\langle J\rangle$ as a function of the ADTS coefficient for currents slightly above threshold. The lines highlight the $1/x^{1/2}$ (bunch size and COM) respective $1/x$ ($\langle J\rangle$ dependency with the dotted line being the mirror of the solid line at negative ADTS.
@@
Figure 12: Measurements of the maximal center-of-mass oscillation amplitude (95percentile) as a function of the ADTS coefficient. The Measurements were taken at bunch currents close to the threshold and with a chromaticity of 0.05. The lines highlight the $1/x^{1/2}$ dependency with the dotted line being the mirror of the solid line at negative ADTS. The errors show the standard deviation between multiple consecutive measurements per point.
@@
Figure 13: Tune shift calculated from the maximal (95percentile) and the minimal (5percentile) action $\langle J\rangle$ as function of ADTS coefficient.
@@
Figure 11 also shows the value of the average action of the particle ensemble $\langle J\rangle$ at the times of minimal bunch size and center-of-mass oscillations [^3] The 5th percentile is taken as value for the minimal action $\langle J\rangle$. as a function of ADTS coefficient, again with the asymmetry for the different signs of the ADTS coefficient visible. The minimal value $\langle J\rangle$ reaches can be connected to the point were the instability is no longer damped and the beam becomes unstable again, as will be described further in the next section (Sec. IV.4).
@@
Both, the maximal bunch size and center-of-mass oscillation amplitudes as well as the minimal action $\langle J\rangle$ show a characteristic dependence on the ADTS coefficient. For $\langle J\rangle$ it follows a $1/x$ dependency and the bunch size as well as the center-of-mass oscillation amplitude has a $1/\sqrt{x}$ dependency. This already hints at a connection with the tune shift via Equ. 3. Going one step further and calculating the tune shift due to ADTS from the simulated values of the minimal action $\langle J\rangle$ for each ADTS coefficient, shows a constant but different level of tune shift for each sign of the ADTS (Fig. 13). While for negative ADTS the calculated tune shift of $\approx 0.00012$ is approximately 8% of the synchrotron tune, the shift for positive ADTS is with $\approx 0.00094$ already 65% of the synchrotron tune. This significant difference gives rise to the hypothesis, that for the positive and negative ADTS a different level of tune shift is required to contain the instability. The tune shift calculated from the maximal $\langle J\rangle$ shows a very similar behavior, where the small difference between maximal and minimal values on the positive side of the ADTS is explained by the fact that nearly no sawtooth is observed for positive ADTS coefficients and currents close to the threshold (compare Fig. 10). Overall, it can be concluded that the tune shift stays between these two levels, with the higher one being the point were the beam stabilises and starts to damp and the lower one indicates when the stabilizing effect stops and the beam goes unstable and blows up again. This is supported in the next section by theoretical calculations evaluating the instability threshold of the transverse mode coupling taking into account the Landau damping.
@@ IV.4 Theoretical Calculations
Refer to caption Refer to caption
Figure 14 shows stability diagrams calculated from Eq. 6 for the cases of positive and negative ADTS. Contours are drawn by plotting the imaginary part of the inverse dispersion relation $I^{-1}_{m}$ against the real part. These contours map out a teardrop shape in complex frequency space pointing towards negative coherent tune shifts in the case of negative ADTS. Changing the sign of the ADTS to positive reflects the contour about the line of zero coherent frequency shift. Also shown on the figure are the eigenvalues of the coupling matrix without Landau damping where two modes are included: the azimuthal head tail modes $m=0$ and $m=-1$. For a head-tail mode to be stable in isolation, its complex coherent frequency shift would have to be within the Landau contour. The condition in the presence of mode coupling is slightly different, given by the zero determinant in Eq. 6, but the images are still illustrative none the less. It can be seen that, in order to influence the stability, a positive ADTS coefficient must be much larger in magnitude than a negative one, matching the observations in measurement and simulation shown in Fig. 11 and 12. This is intuitive as a negative ADTS coefficient means that the tune spread is towards negative tune shifts and the current-dependent tune shift of the $m=0$ mode is also negative for most broadband impedances.
One feature of storage rings used for fourth-generation synchrotron light sources, particularly those using low-RF frequencies (such as the 100 MHz of the 3 GeV ring at MAX IV), that is beneficial in this regard is the low incoherent synchrotron frequency. This means that the coupling frequency of the $m=0$ and $m=-1$ head-tail modes is not so far out of the spread in betatron tune of the electron bunches. Nevertheless, both in simulation and measurement at the 3 GeV ring at MAX IV, the magnitude and sign of the ADTS coefficient does not impact the threshold current of the TMCI. As discussed in Sec. IV.1, this is because the point of the mode-coupling is outside the tune spread of the bunch when it is stable. When the bunch goes unstable and increases in size, however, the tune spread increases until the Landau damping kicks in. Figure 15
Figure 15: Predicted equilibrium actions $\langle J\rangle$ and those determined from simulations in Mbtrack2 for an ADTS coefficient $b=1000\,\mathrm{tm^{-1}}$.
shows the saturation equilibrium average particle actions $\langle J\rangle$ in simulation and from the results of Eq. 6 for the two azimuthal head-tail modes and for a given ADTS coefficient $b$. Radiation damping is included in the simulations so they show the same sawtooth behavior seen in the measurements. The saturation action $\langle J\rangle$ is then taken as the minimum in this sawtooth pattern, corresponding to the point where the damping stops and the bunch blows up again, because during the rest of the sawtooth period, the beam is either being damped (and is therefore stable) or has some centroid motion, which has a negative effect on the Landau damping (this is a potential root cause of the sawtooth behavior). The agreement is not perfect, particularly in the value of the threshold current. This may change with the number of head-tail modes included in the theoretical prediction, although including more significantly complicates the numerical optimisation.
IV.5 Betatron Tune Shift with Current
@@
Figure 16: Coherent motion spectrum showing the current dependent betatron tune shift below and above the instability threshold. Simulation (top): Fourier transform of the center-of-mass oscillation plotted as a function of the bunch current for negative (left) and positive (right) ADTS coefficient $b=15000$/m. Measurement(bottom): Fourier transform of the center-of-mass oscillation as a function of bunch current for an ADTS coefficient of $b=-10000$/m (left) and $b=13720$/m. During the measurement the instability was “switched” on and off (see text).
@@
Observing the vertical betatron tune as function of current directly shows the expected current-dependent tune shift due to the transverse impedance from the zero-current tune of $\nu_{0}=0.275$ towards the -1 mode at the first synchrotron frequency side band ($\nu_{0}-0.00146$). In simulations of the coherent beam spectrum the threshold of the TMCI is clearly visible as the current at which the tune couples to the -1 mode (top row in Fig. 16). This is the same for both signs of the ADTS. The difference for negative and positive ADTS starts above the threshold where for negative ADTS the tune seems to continue its shift towards a lower tune with a similar slope as below the threshold (Fig. 16a). For the positive ADTS, the behaviour looks very different. While a slight shift in the opposite direction to higher tunes would not be unexpected, due to the positive sign of the ADTS, the tune jumps within a very small current range above the threshold from the -1 mode back to the 0 mode and then shows a continuous shift to higher tunes from there (Fig. 16b).
@@
This drastic difference in behaviour can also be seen in measurements. Figures 16c and 16d show the measured tune spectra at different bunch currents and for negative and positive ADTS respectively. The measurements were conducted in such a way that the previously described hysteresis of the instability threshold (Sec. IV.1) was used to get comparative measurements for the tune of a stable and an unstable beam. To this end, the measurement was started at high bunch currents and the tune spectrum was recorded alternately for a beam stabilized by the BBB feedback system [^4] After initial stabilization the feedback is switched off during the measurement. and for an unstable beam, where the instability was triggered by a short excitation [^5] The excitation is switched off as well before the measurement is taken..
@@
For the stable beam, the tune continues its current-dependent shift towards lower values. For both signs of the ADTS, it is clearly visible that the presence of the instability shifts the tune compared to the tune of the stable beam. For negative ADTS (Fig. 16c), the shift is small and towards slightly lower tune values. For positive ADTS (Fig. 16d), the tune is shifted back towards the zero-current tune (0 mode) and shows a very small current-dependent shift towards higher tune. Except for the difference in threshold and the threshold hysteresis observed in the measurements, the simulation and the measurements agree very well with respect to the tune shifts below and above threshold.
@@
At higher bunch currents, additional features appear. In the measurement at negative ADTS, an upper and lower sideband shows up, moving with the tune as function of current. In the case of the positive ADTS the tune peak is broadened greatly and nearly spans from the -1 mode to the +1 mode. Comparing with the calculated tune shift of $\approx 0.001$ resulting from the maximal action $J$ simulated in case of positive ADTS (Fig. 13), shows that the jump of the coherent betatron tune by one synchrotron tune ($\nu_{s}=0.00146$) towards the 0 mode is only slightly bigger. Overall, from the measurement and tracking simulations it is not apparent whether this difference in the behaviour of the coherent tune above threshold is the cause or a consequence of the observed asymmetry in the level of beam “blow up” for negative versus positive ADTS.
@@
Landau damping has been investigated in the past as a possible source as mitigation mechanism of mode-coupling instabilities, also in connection with the amplitude-dependent tune shift as the source of the required tune spread.
@@
While during standard operations the bunch current at the 3 GeV ring at the MAX IV Laboratory is below the TMCI threshold, an asymmetric dependence on the sign of the amplitude-dependent tune shift (ADTS) has been previously observed in dedicated experiments. Systematic studies were now conducted to investigate this observed asymmetry in dedicated single bunch experiments. It was observed, that for some ADTS coefficients the beam was lost when crossing the threshold while at others a saw-tooth shaped amplitude modulation was observed on the center-of-mass oscillation as well as on the bunch size leading to a self-contained instability. The presented simulations with the tracking tool mbtrack2 and the conducted measurements are in good agreement. Both show that the observed threshold current is independent of the ADTS coefficient and an observed hysteresis in the measured threshold can be attributed to intra-beam scattering effects. For the dynamics above the threshold, both measurements and simulation, show that for positive ADTS coefficients the maximal center-of-mass oscillation amplitude and bunch size that is reached before the instability stabilizes is systematically higher than for negative ADTS coefficients, indicating that this could be the cause of the observed partial beam current losses. The same asymmetry is also visible in the tune shift at the minimal $\langle J\rangle$ required for damping. The shift is constant and the value is only dependent on the sign of the ADTS, with the tune shift calculated for positive ADTS coefficients already being at $\approx 65\%$ of the synchrotron tune. Stability diagrams with Landau contours, calculated to include the amplitude-dependent tune shift, show as well that higher positive ADTS coefficients compared to negative ones are required for the instability to be Landau damped. Furthermore, simulations and measurements of the coherent tunes as a function of the bunch current, show a strong difference in the tunes development above threshold. For negative ADTS coefficients the tune is slightly shifting to lower values starting from the -1 mode at the threshold. In contrast, for positive ADTS coefficients, the coherent tune jumps back to the 0 mode and only then shows with increasing current a slight shift to higher tune values, as might be expected for positive ADTS coefficients.

### v2

@@ IV.2 Beam Losses at Threshold
A significant influence of the magnitude and sign of the ADTS coefficients can be observed in the amount of charge lost when the instability threshold is crossed during injection. Figure 7 shows the beam loss in percent for different values of the ADTS coefficient with a chromaticity close to zero or with slightly negative chromaticity values. For negative ADTS coefficients up to nearly zero ($\approx-500/\mathrm{m}$) no beam loss is encountered at all when crossing the instability threshold during injection. This is already noteworthy as it shows that the instability is not destructive even though it leads to strong center-of-mass oscillations and an increase in bunch size. On the other side, for positive ADTS coefficients a partial beam loss is observed when crossing the threshold. For values from zero up to $6000/\mathrm{m}$ more than 40 and up to 90 percent of the beam current is lost. For higher positive ADTS coefficients, the loss goes down close to zero again. So, for ADTS coefficients up to $6000/\mathrm{m}$, there is a difference in the observed behaviour for a positive and negative sign of the ADTS coefficient. While at negative coefficients the instability is self-containing, for positive coefficients a partial beam loss is observed until the beam stabilizes again.
@@ IV.3 Bunch Position and Size
Figure 8: Synchronized measurement of vertical spot profile and center-of-mass amplitude as function of time. The image shows the vertical light spot measured at different points in the sawtooth like dynamic. In red the envelope of the center-of-mass motion is displayed with a point for each spot profile measurement. The yellow dots indicate the spot size (std) gained from a Gaussian fit. Measured at ADTS coefficient $b=-10000/\mathrm{m}$ and current of 2.6 mA.
For a negative ADTS coefficient, the dynamic above threshold shows clear, regular, sawtooth like bursts in transverse bunch size and as amplitude modulation of the center-of-mass oscillations. This is visible in the BPM trace directly (Fig. 2) as well as in the synchronous measurement of bunch position and bunch size in Fig. 8. The contribution of the bunch size can be seen in the difference in behavior over turns between the vertical beam-size and centre of mass. Additionally, the calculated spot size of up to 0.68 mm (standard deviation determined by Gaussian fit) in Fig. 8 compared to the detected center-of-mass oscillation amplitude of maximal 0.4 mm (= 0.8 mm peak-peak) indicates that the bunch size has a non-negligible contribution. Furthermore, the center-of-mass oscillation goes back to nearly zero for some hundred turns between the increases in oscillation amplitude. During this time the observed spot size slowly damps down indicating that the bunch size is damping down, reaching minimal values in the order of 0.35 mm before the next peak. The same sawtooth behaviour is present in the simulation shown in Fig. 9 when synchrotron radiation effects, such synchrotron radiation losses, synchrotron radiation damping and quantum excitations are included.
Figure 9: Simulated center-of-mass oscillation and bunch size at an ADTS coefficient $b=-15000/\mathrm{m}$ and a current of 4.8 mA.
Refer to captionRefer to captionRefer to captionRefer to caption
@@
This dynamic indicates a stabilizing mechanism which leads to a containment of the instability instead of a continuous growth until charge is lost. The behavior in the amplitude of the center-of-mass oscillations and the bunch size show that at one point a temporary stabilization occurs which leads to a damping of the oscillation to below the noise limit of the measurement. The bunch size is also damped down during this stable period but it does not reach the expected stable bunch size before the instability is triggered again leading to a fast blow-up of the bunch size and the onset of strong center-of-mass oscillations. A possible mechanism behind these dynamics, is Landau damping. The blown-up bunch size, increased COM oscillations and the ADTS lead to an increased tune spread which can cause a temporally increased Landau damping effect. The result can be described as a “self-containing” instability, caused by Landau damping, which only sets in when the bunch is blown-up and the ADTS results in a significant tune shift. Two points can be identified in the dynamics. An “upper turning point” when the damping due to the ADTS caused by the blow-up of the beam becomes predominant and overpowers the growth driven by the instability. And a “lower turning point” when the bunch size and COM oscillations are damped down to such low values that the resulting ADTS is not enough to Landau damp the instability any longer and the growth starts again.
While the IBS is considered as the cause of the hysteresis observed in the Instability threshold, it is not a candidate for explaining the self-containing dynamics, as it operates in the wrong sense, ie. a blown-up beam has a lower threshold current and is therefore more unstable and does not contribute to a self-containing effect, where the threshold would need to increase to temporarily stabilize the unstable beam.
@@
For positive ADTS, the dynamics above the threshold can only be observed in measurements at high ADTS coefficients where no instantaneous charge loss occurs. For higher bunch currents, the dynamics in the center-of-mass oscillation and the bunch size have a similar sawtooth like pattern as observed for negative ADTS coefficients (see left side of Fig. 10). The pattern changes for lower bunch currents, closer to the threshold, as shown on the right side in Fig. 10. Here, the center-of-mass oscillation amplitude is only lightly modulated in the measurements, and nearly constant in the simulations.
@@
The simulation results show that a positive ADTS (Fig. 10, lower left plot) leads to a stronger blow-up than a negative ADTS (Fig. 9). The experiments (Fig. 2 and Fig. 10, upper left plot) show the same asymmetric behaviour This indicates, that the level of blow-up at which the instability is contained (the upper turning point) and finds some kind of equilibrium, pseudo-stable state is different for negative and positive ADTS. The asymmetry is clearly visible in Fig. 11 where, for a bunch current slightly above threshold, the maximal bunch size and the maximal oscillation amplitude of the center-of-mass is given as a function of ADTS coefficient in simulations [^2] To be more robust against outliers the 95th percentile of the bunch size and the center-of-mass oscillation amplitude are taken.. The range in ADTS coefficient where partial current loss occurs depends on the combination of the center-of-mass oscillations and the total bunch size which, above a certain value, leads to parts of the charge being “scraped” by the beam pipe. Independent of the exact value, it can be seen from Fig. 11 that the affected ADTS range would not be symmetric around zero but rather shifted to positive ADTS coefficients. The measurements show the same dependence of the maximal center-of-mass oscillation amplitude [^3] Again, the 95th percentile of center-of-mass oscillation is taken. as a function of the ADTS coefficient (Fig. 12) for negative ADTS. While it is not measurable at the lower positive ADTS coefficients, due to the partial beam losses, the measured values at higher positive ADTS coefficients are higher than the corresponding values at negative ADTS showing the same asymmetry as in the simulations in Fig. 11. So to reach the same level of suppression of the instability, meaning low values in maximal bunch size and center-of-mass oscillations, a higher positive than negative ADTS coefficient is be needed. This asymmetry observed in both measurement and simulations provides the experimental explanation for the asymmetry observed in the beam loss at the threshold (see Fig. 7).
@@
Figure 11: Simulated maximal (95th percentile) bunch size and center-of-mass oscillation amplitude and minimal (5th percentile) action $\langle J\rangle$ as a function of the ADTS coefficient for currents slightly above threshold. The lines highlight the $1/x^{1/2}$ (bunch size and COM) respective $1/x$ ($\langle J\rangle$) dependency with the dotted line being the mirror of the solid line at negative ADTS.
@@
Figure 12: Measurements of the maximal center-of-mass oscillation amplitude (95th percentile) as a function of the ADTS coefficient. The measurements were taken at bunch currents close to the threshold and with a chromaticity of 0.05. The lines highlight the $1/x^{1/2}$ dependency with the dotted line being the mirror of the solid line at negative ADTS. The errors show the standard deviation between multiple consecutive measurements per point.
@@
Figure 13: Tune shift calculated from the maximal (95th percentile) and the minimal (5th percentile) action $\langle J\rangle$ as function of ADTS coefficient.
@@
Figure 11 also shows the value of the average action of the particle ensemble $\langle J\rangle$ at the times of minimal bunch size and center-of-mass oscillations [^4] The 5th percentile is taken as value for the minimal action $\langle J\rangle$. as a function of ADTS coefficient, again the asymmetry for the different signs of the ADTS coefficient is visible. The minimal value $\langle J\rangle$ reaches can be connected to the point were the instability is no longer damped and the beam becomes unstable again, the lower turning point of the dynamics.
(a) Negative ADTS
(b) Positive ADTS
@@
Both, the maximal bunch size and center-of-mass oscillation amplitudes as well as the minimal action $\langle J\rangle$ show the characteristic dependence on the ADTS coefficient. For $\langle J\rangle$ it follows a $1/x$ dependency and the bunch size as well as the center-of-mass oscillation amplitude has a $1/\sqrt{x}$ dependency. This again indicates the connection with the tune shift, via Eq. 3. Going one step further and calculating the tune shift due to ADTS from the simulated values of the minimal action $\langle J\rangle$ for each ADTS coefficient, shows a constant but different level of tune shift for each sign of the ADTS (Fig. 13). While for negative ADTS the calculated tune shift of $\approx 0.00012$ is approximately 8% of the synchrotron tune, the shift for positive ADTS is with $\approx 0.00094$ already 65% of the synchrotron tune. This significant difference implies, that for positive and negative ADTS a different level of tune shift is required to contain the instability via sufficient Landau damping. The tune shift calculated from the maximal $\langle J\rangle$ shows a very similar behavior, where the small difference between maximal and minimal values on the positive side of the ADTS is explained by the fact that nearly no sawtooth is observed for positive ADTS coefficients and currents close to the threshold (compare Fig. 10). We conclude that the tune shift stays between these two levels, with the higher one being the point were the beam stabilises and starts to damp (upper turning point) and the lower one indicates when the stabilizing effect stops and the beam goes unstable and blows up again (lower turning point). And while the overall behaviour is the same for both signs of the ADTS, the tune shift at the turning points is significantly higher for positive ADTS coefficients.
That this can be attributed to Landau damping, is supported in the following section by theoretical calculations comparing the instability growth rates of the transverse mode coupling against Landau contours.
@@ IV.4 Theoretical Calculations
Figure 14 shows stability diagrams calculated from Eq. 4 for the cases of positive and negative ADTS. The calculations were performed with an internal code which was bench marked against DELPHI [17]. Contours are drawn by plotting the imaginary part of the inverse dispersion relation $I^{-1}_{m}$ against the real part. These contours map out an asymmetric shape in complex frequency space pointing towards negative coherent tune shifts in the case of negative ADTS. Changing the sign of the ADTS to positive reflects the contour about the line of zero coherent frequency shift. Also shown on the figure are the eigenvalues of the coupling matrix in Eq. 6 without Landau damping where three modes are included: the azimuthal head tail modes $m=0$, $m=-1$ and $m=-2$. As radial modes have been neglected in order to make the image clearer, it can be observed (contrary to measurements) that the beam becomes stable again at higher currents. For a head-tail mode to be stable in isolation, its complex coherent frequency shift would have to be within the Landau contour. The condition in the presence of mode coupling is slightly different and given by the zero determinant in Eq. 6. How these diagrams are best applied to the TMCI to solve Eq. 6 is an active research topic, discussed for example in [2]. Nonetheless, the images in Fig. 14 are still illustrative. It can be seen that, in order to influence the stability, a positive ADTS coefficient must be much larger in magnitude than a negative one, matching the observations in measurement and simulation shown in Fig. 11 and 12. This is intuitive as a negative ADTS coefficient means that the tune spread is towards negative tune shifts and the current-dependent tune shift of the $m=0$ mode is also negative for most broadband impedances. Therefore, the shifted coherent tune frequency lies within the band of the incoherent tune and motions with this tune are damped by the Landau damping effect.
@@
Figure 15: Coherent motion spectrum showing the current-dependent betatron tune shift below and above the instability threshold. Simulation (top): Fourier transform of the center-of-mass oscillation plotted as a function of the bunch current for negative (left) and positive (right) ADTS coefficient $b=15000$/m. Measurement(bottom): Fourier transform of the center-of-mass oscillation as a function of bunch current for an ADTS coefficient of $b=-10000$/m (left) and $b=13720$/m. During the measurement the instability was “switched” on and off.
One feature of storage rings used for fourth-generation synchrotron light sources, particularly those using low-RF frequencies (such as the 100 MHz of the 3 GeV ring at MAX IV), that is beneficial in this regard is the low incoherent synchrotron frequency. This means that the coupling frequency of the $m=0$ and $m=-1$ head-tail modes is not so far out of the spread in betatron tune of the electron bunches. Nevertheless, both in simulation and measurement at the 3 GeV ring at MAX IV, the magnitude and sign of the ADTS coefficient does not impact the threshold current of the TMCI. As discussed in Sec. IV.1, this is because the point of the mode-coupling is outside the tune spread of the bunch when it is stable. When the bunch goes unstable and increases in size, however, the tune spread increases until the Landau damping kicks in.
IV.5 Betatron Tune Shift with Current
@@
Observing the vertical betatron tune as function of current directly shows the expected current-dependent tune shift due to the transverse impedance from the zero-current tune of $\nu_{0}=0.275$ towards the -1 mode at the first synchrotron frequency side band ($\nu_{0}-0.00146$). In simulations of the coherent beam spectrum the threshold of the TMCI is clearly visible as the current at which the tune couples to the -1 mode (top row in Fig. 15). This is the same for both signs of the ADTS. The difference for negative and positive ADTS starts above the threshold current where for negative ADTS the tune continues its shift towards a lower tune with a similar slope as below the threshold (Fig. 15a). For the positive ADTS, the behaviour looks very different. While a slight shift in the opposite direction to higher tunes would not be unexpected, due to the positive sign of the ADTS, the tune jumps within a very small current range above the threshold from the -1 mode back to the 0 mode and then shows a continuous shift to higher tunes from there (Fig. 15b).
@@
This drastic difference in behaviour can also be seen in measurements. Figures 15c and 15d show the measured tune spectra at different bunch currents and for negative and positive ADTS, respectively. The measurements were conducted in such a way that the previously described hysteresis of the instability threshold (Sec. IV.1) was used to get comparative measurements for the tune of a stable and an unstable beam. To this end, the measurement was started at high bunch currents and the tune spectrum was recorded alternately for a beam stabilized by the BBB feedback system [^5] After initial stabilization the feedback is switched off during the measurement. and for an unstable beam, where the instability was triggered by a short excitation [^6] The excitation is switched off as well before the measurement is taken..
@@
For the stable beam, the tune continues its current-dependent shift towards lower values. For both signs of the ADTS, it is clearly visible that the presence of the instability shifts the tune compared to the tune of the stable beam. For negative ADTS (Fig. 15c), the shift is small and towards slightly lower tune values. For positive ADTS (Fig. 15d), the tune is shifted back towards the zero-current tune (0 mode) and shows a very small current-dependent shift towards higher tune. Except for the difference in threshold and the threshold hysteresis observed in the measurements, the simulation and the measurements agree qualitatively very well with respect to the tune shifts below and above threshold.
@@
At higher bunch currents, additional features appear. In the measurement at negative ADTS, an upper and lower sideband shows up, moving with the tune as function of current. In the case of the positive ADTS the tune peak is broadened greatly and nearly spans from the -1 mode to the +1 mode. Comparing with the calculated tune shift of $\approx 0.001$ resulting from the maximal action $J$ simulated in case of positive ADTS (Fig. 13), shows that the jump of the coherent betatron tune by one synchrotron tune ($\nu_{s}=0.00146$) towards the 0 mode is only slightly bigger. Overall, from the measurement and tracking simulations it is not apparent whether this difference in the behaviour of the coherent tune above threshold is the cause or a consequence of the observed asymmetry in the level of beam blow-up for negative versus positive ADTS.
@@
Landau damping has been investigated in the past as a possible mitigation mechanism of mode-coupling instabilities, also in connection with the amplitude-dependent tune shift as the source of the required tune spread, mainly at proton machines. In this paper, we showed how Landau damping can affect the dynamics of the transverse mode-coupling instability under certain operational parameters at a fourth generation synchrotron light source. While during routine operations the bunch current at the 3 GeV ring at the MAX IV Laboratory is below the TMCI threshold, an asymmetric dependence on the sign of the ADTS has been previously observed in dedicated experiments. Systematic studies were now conducted to investigate this observed asymmetry in dedicated single bunch experiments. It was observed, that for some ADTS coefficients the beam was lost when crossing the threshold while at others a saw-tooth shaped amplitude modulation was observed on the center-of-mass oscillation as well as on the bunch size leading to a self-contained instability.
@@
The presented simulations with the tracking tool mbtrack2 and the conducted measurements are in good agreement. Both show that the observed threshold current is independent of the ADTS coefficient and an observed hysteresis in the measured threshold can be attributed to intrabeam scattering effects. For the dynamics above the threshold, both measurements and simulation, show that for positive ADTS coefficients the maximal center-of-mass oscillation amplitude and bunch size, that is reached before the instability stabilizes, is systematically higher than for negative ADTS coefficients, indicating that this could be the cause of the observed partial beam current losses. The same asymmetry is also visible in the tune shift at the minimal $\langle J\rangle$ required for damping. The shift is constant and the value is only dependent on the sign of the ADTS, with the tune shift calculated for positive ADTS coefficients already being at $\approx 65\%$ of the synchrotron tune. Stability diagrams with Landau contours, calculated to include the amplitude-dependent tune shift as well show that higher positive ADTS coefficients compared to negative ones are required for the instability to be Landau damped. Furthermore, simulations and measurements of the coherent tunes as a function of the bunch current, show a strong difference in the tunes development above threshold. For negative ADTS coefficients the tune is slightly shifting to lower values starting from the -1 mode at the threshold. In contrast, for positive ADTS coefficients, the coherent tune jumps back to the 0 mode and only then shows with increasing current a slight shift to higher tune values, as might be expected for positive ADTS coefficients.